# Optimizing Visible Light Communication Efficiency Through Reinforcement Learning-Based NOMA-CSK Integration

### v1

@@ 4 Simulations
[…]
@@
Here, $S_{i}^{(1)}$ and $S_{i}^{(2)}$ are the value of the CSK symbol in color $i$ for the $U_{1}$ and $U_{2}$, respectively. $\rho^{(1)}$ is the NOMA power allocation coefficient for $U_{1}$, who has a bigger channel gain. ${{h^{(k)}}}$ is the total channel between tranmister and $k-th$ user. In this case, the SINR at the receivers when demodulation is performed with the help of SIC is:
$\operatorname{SINR}^{(1)}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}^{(1)}\delta_{ij}}{\sum\limits_{i}\sum\limits_{j}{({p_{i,j}}-p_{i,j}^{(1)}\delta_{ij})}+N_{0}^{(1)}}$ (11)
$\operatorname{SINR}^{(2)}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}^{(2)}\delta_{ij}}{\sum\limits_{i}\sum\limits_{j}{({p_{i,j}}-p_{i,j}^{(2)}\delta_{ij})}+N_{0}^{(2)}+N^{(1)}}$ (12)
$N_{0}^{(k)}=\frac{\left(\sum\limits_{i}\sum\limits_{j}\left|p_{i,j}^{(k)}\delta_{ij}\right|\right)^{2}}{SNR}$ (13)
@@
$N_{0}^{(1)}$ and $N_{0}^{(2)}$ represent the AWGN noise powers for $U_{1}$ and $U_{2}$ respectively. $N^{(1)}$ denotes the interference caused by the message signals of $U_{1}$ on $U_{2}$, and it can be expressed as follows:
$N^{(1)}=\sum_{i}\sum_{j}\left(\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}S_{i}^{(1)^{2}}p^{(1)}\left(h^{(1)}\right)^{2}\mathrm{~{}d}\lambda\right)$ (14)
@@
In a practical communication system, the positions of users are not fixed, and the channel conditions change as they move. These factors have been considered in the proposed reinforcement learning-based NOMA-CSK power allocation method. Our objective was to train the agent to efficiently allocate power even when confronted with channel conditions that differ from those encountered during training. Reinforcement learning is a method for training an agent in a simulation environment through trial and error, without relying on a pre-existing dataset. During these trials, the agent learns a policy that maximizes rewards by receiving a positive reward when it achieves a desired outcome, and a negative reward (penalty) otherwise. In this study, a model-free, online, off-policy, actor-critic reinforcement learning agent called soft actor critic (SACA) was utilized. Since both the action and observation values are continuous, an agent capable of working in this environment is selected as the SACA [27].
The environment should be designed before the learning process can be implemented. The actions that the agent can take and the quantities it can observe should be determined to design the environment. The action to be taken by the agent is defined as changing the power allocation $\rho$ of the user with the highest channel gain within the range of [0 0.5]. In this case, the remaining power will be used by other users. The value that the agent will observe is the ratio of the channel gains of $U_{1}$ and $U_{2}$ to each other is given in Eq.15, and it is randomly determined.
$\mathrm{r}=\frac{h^{(2)}}{h^{(1)}},\mathrm{r}\in\left[\begin{array}[]{ll}0&1\end{array}\right]$ (15)
@@
The users’ channel gains are determined to fall within the range of $2.84\times 10^{-5}$ and $5.98\times 10^{-4}$. This range encompasses the maximum and minimum values among 2601 channels, obtained by varying the users’ coordinates as $4.50.5<X<$ 4.5, $4.50.5<Y<$ 4.5, and $2.50.5<Z<$ 2.5, with an increment of 0.25 for each step. The values of $h^{(1)}$ and $h^{(2)}$ are generated randomly according to Eq.16 and Eq.17, respectively.
@@
After creating the environment, the agent’s performance criteria and reward mechanism should be determined. The agent will be trained to maximize fairness as an important objective in multiple access systems. Fairness is a measure of how fairly a system resource is shared among users or applications used in communication systems and is given in Eq.18.
$\frac{\left(\sum\limits_{k=1}^{N}T_{k}\right)^{2}}{N\sum\limits_{k=1}^{N}T_{k}^{2}}$ (18)
@@
Here, $N$ represents the total number of users in the system. $T_{k}$ represents the throughput of the $k-th$ user. If the fairness value that can be calculated using Eq.18 is above 0.99, the agent is rewarded with 1; otherwise, it is rewarded with -10. The SACAgent is trained for 1000 episodes by randomly generating user channel ratios and channels. Training is completed when the average reward value of the last 20 episodes is 1. The agent is tested with 100 randomly generated pairs of $({{h^{(1)}}},{{h^{(2)}}})$ where the value of $r$ is varied in increments of 0.01 within the range of 0 to 1.
@@
The changes in sum-rate, throughput, and fairness values are depicted in Fig.1(a), Fig.1(b), and Fig.1(c) as the channels ratio $r$ varies. These values were obtained by applying proposed power allocation techniques to users in randomly generated $({{h^{(1)}}},{{h^{(2)}}})$ channels. The figures also include the performance results of TDMA-based multiple access, GRPA and NGDPA. Upon examination, it becomes clear that the trained agent demonstrates significantly superior performance in sum-rate, throughput, and fairness compared to other methods.
@@
Figure 2: Performance comparison of the agent, grpa, ngdpa and tdma
For a 30 dB SNR, a bit error rate (BER) of 0 was achieved for all channel rates in TDMA and AGENT, while the BER graphs obtained for GRPA and NGDPA are shown in Fig.1(d).
The PSD of the LED’s light output is altered due to modulation. In order to determine CRI , CCT, and illumination level, the modified PSD is used for calculations. Eq. 19 represents the power of the transmitted symbols of $U_{1}$ and $U_{2}$ in color $i$. By this, altered PSD ($PSD_{a}(\lambda)$) can be calculated using Eq. 20
$TSP_{i}=\mathbb{E}\left(\left[S_{i}^{(1)}\sqrt{p^{(1)}}+S_{i}^{(2)}\sqrt{1-p^{(1)}}\right]^{2}\right)$ (19)
$PSD_{a}(\lambda)=\sum\limits_{i=1}^{3}PSD_{i}(\lambda)TSP_{i}$ (20)
In this case, after obtaining $PSD_{a}$, CRI, CCT and illumination parameters for the cases with and without communication are calculated and given in Table 3. The NOMA method did not change the CRI, CCT and illumination level. Only a decrease in illumination value was observed due to modulation in communication.
Table 3: Illuminations parameters comparision table
CRI CCT Illumination (lux)
W/O Comm. 32.3121 9445 863.64
W Comm. 32.3121 9445 287.59
The agent was trained for a specific channel range. To investigate its performance in an environment other than where it was trained, we tested it in a new environment with a lower bound value 5 times smaller and an upper bound value 5 times larger than that of the environment it was trained in. We compared the agent’s performance with other power allocation methods. As can be seen from Fig.2(a), Fig.2(b), and Fig.2(c), the agent performs better in terms of sum-rate, throughput, and fairness compared to other methods in an environment outside of where it was trained. The BER graph for NGDPA and GRPA is given in Fig.2(d) while 0 BER is achieved with AGENT and TDMA.

### v2

@@ 4 Numerical Results
$p_{i,j}^{(k)}=\frac{1}{3}\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}(\lambda)\left[S_{i}^{(1)}{\rho^{(1)}}+S_{i}^{(2)}{(1-\rho^{(1)})}\right]h^{(k)}d\lambda$ (10)
@@
Here, $S_{i}^{(1)}$ and $S_{i}^{(2)}$ are the value of the CSK symbol in color $i$ for $U_{1}$ and $U_{2}$, respectively. $\rho^{(1)}$ is the NOMA power allocation coefficient for $U_{1}$, who has a better channel gain ($h^{(1)}>h^{(2)}$). The total channel gain between the transmitter and the $k$-th user is $h^{(k)}$.
The received power component at user $k$ originating from the signal intended for user $u\in\{1,2\}$ is given by:
$P_{i,j}^{(k,u)}=\frac{1}{3}\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}(\lambda)S_{i}^{(u)}\rho^{(u)}h^{(k)}d\lambda$ (11)
where $\rho^{(2)}=1-\rho^{(1)}$.
The total desired signal power for user 1 ($U_{1}$) and user 2 ($U_{2}$) are:
$S^{(1)}=\sum_{i,j}P_{i,j}^{(1,1)}\delta_{ij}\quad\text{and}\quad S^{(2)}=\sum_{i,j}P_{i,j}^{(2,2)}\delta_{ij}$ (12)
The cross-color interference for user $k$ from its own signal is $I_{CCI}^{(k)}=\sum_{i,j}P_{i,j}^{(k,k)}(1-\delta_{ij})$.
With Successive Interference Cancellation (SIC), user 1 first decodes and removes user 2’s signal. However, since user 2 has higher power, its signal acts as interference for user 1. Conversely, user 2 sees user 1’s signal as interference. The inter-user interference experienced by user 2 is $I^{(2\leftarrow 1)}=\sum_{i,j}P_{i,j}^{(2,1)}$.
The SINR at the receivers can now be expressed unambiguously:
$\operatorname{SINR}^{(1)}=\frac{S^{(1)}}{I_{CCI}^{(1)}+\sum_{i,j}P_{i,j}^{(1,2)}+N_{0}^{(1)}}$ (13)
$\operatorname{SINR}^{(2)}=\frac{S^{(2)}}{I_{CCI}^{(2)}+I^{(2\leftarrow 1)}+N_{0}^{(2)}}$ (14)
@@
$N_{0}^{(1)}$ and $N_{0}^{(2)}$ represent the AWGN noise powers for $U_{1}$ and $U_{2}$ respectively. $I^{(1)}$ denotes the interference caused by the message signals of $U_{1}$ on $U_{2}$, and it can be expressed as follows:
$I^{(1)}=\sum_{i}\sum_{j}\left(\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}S_{i}^{(1)}\rho^{(1)}h^{(1)}\mathrm{~d}\lambda\right)$ (15)
@@
In a practical communication system, the positions of users are not fixed, and the channel conditions change as they move. These factors have been considered in the proposed reinforcement learning-based NOMA-CSK power allocation method. Our objective was to train the agent to efficiently allocate power even when confronted with channel conditions that differ from those encountered during training.
To provide a clear baseline for comparison, we detail the established power allocation schemes used as benchmarks. It is crucial to reiterate that these methods were developed for scenarios without the complex cross-color interference present in our NOMA-CSK model, but we implement them here to establish a performance benchmark.
GRPA-NOMA: This method allocates power proportionally to the square of the channel gains between users. The powers allocated to users 1 and 2 ($P_{1},P_{2}$) are determined as follows, assuming user 1 has the better channel condition ($h^{(1)}>h^{(2)}$):
${P_{1}=\frac{P_{total}}{1+\frac{(h^{(1)})^{2}}{(h^{(2)})^{2}}}}$ (16)
$P_{2}=P_{total}-P_{1}$ (17)
Here, $P_{total}$ is the total transmission power. The power allocation coefficient for user 1 is $\rho^{(1)}=P_{1}/P_{total}$.
NGDPA-NOMA: This method considers the normalized difference between channel gains to determine power allocation. The power allocation is given by:
$P_{1}=\frac{P_{total}}{1+\frac{h^{(1)}}{|h^{(1)}-h^{(2)}|}}$ (18)
$P_{2}=P_{total}-P_{1}$ (19)
For both methods, the resulting throughput for each user is calculated based on their respective SINR values as defined in Eq. 13 and 14.
Reinforcement Learning: Reinforcement learning is a method for training an agent in a simulation environment through trial and error, without relying on a pre-existing dataset. During these trials, the agent learns a policy that maximizes rewards by receiving a positive reward when it achieves a desired outcome, and a negative reward (penalty) otherwise. In this study, a model-free, online, off-policy, actor-critic reinforcement learning agent SAC is utilized. Since both the action and observation values are continuous, an agent capable of working in this environment is selected as the SAC agent [33].
The environment should be designed before the learning process can be implemented. The actions that the agent can take and the quantities it can observe should be determined to design the environment.
To formulate the problem within the RL framework, we define the state, action, and reward function.
  • State (Observation) Space ($\mathcal{S}$): The agent observes the state of the channel, which is captured by the ratio of the users’ channel gains. The state $s_{t}$ at time step $t$ is:
    $s_{t}=\frac{h^{(2)}}{h^{(1)}}\in[0,1]$ (20)
    where $h^{(1)}$ and $h^{(2)}$ are the channel gains of the stronger and weaker users, respectively.
  • Action Space ($\mathcal{A}$): The agent’s action, $a_{t}$, is to select the NOMA power allocation coefficient $\rho^{(1)}$ for the user with the better channel gain. To ensure the weaker user receives more power, this coefficient is bounded.
    $a_{t}=\rho^{(1)}\in[0,0.5]$ (21)
    The power for the second user is then $1-\rho^{(1)}$.
  • Reward Function ($R_{t}$): The reward function guides the agent’s learning process. It is a multi-objective function designed to maximize fairness and overall performance, as detailed in Eq. 24.
@@
The users’ channel gains are determined to fall within the range of $2.84\times 10^{-5}$ and $5.98\times 10^{-4}$. This range encompasses the maximum and minimum values among 2601 channels, obtained by varying the users’ coordinates as $0.5<X<4.5$, $0.5<Y<4.5$, and $0.5<Z<2.5$, with an increment of 0.25 for each step. The values of $h^{(1)}$ and $h^{(2)}$ are generated randomly according to Eq.22 and Eq.23, respectively.
@@
After exploring the combination of NOMA and CSK for VLC systems, the subsequent critical step involves delineating the agent’s performance criteria and configuring its reward mechanism. This pivotal phase is essential for guiding the training regimen, which is facilitated by the SAC algorithm, as delineated in Algorithm 1. SAC is renowned for its effectiveness in addressing complex decision-making tasks.
Central to the SAC algorithm’s efficacy is the precise calibration of various parameters, detailed in Table 3. The hyperparameters in Table 3 were selected based on common practices in DRL literature and refined through preliminary experiments.
Table 3: RL Parameters
Parameter Value
Fairness observation space $O$ $[0,1]$
Power allocation action space AA $[0,0.5]$
Discount factor $\gamma$ 0.99
Entropy coefficient $\alpha$ 1
Target smoothing coefficient $\tau$ $10^{-3}$
Learning rate $\eta$ $3\times 10^{-4}$
Replay buffer size $B$ $10^{4}$
Mini-batch size $M$ 64
MaxEpisodes 1000
MaxStepsPerEpisode 500
StopTrainingCriteria AverageReward
ScoreAveragingWindowLength 100
StopTrainingValue 163
The agent’s training is guided by a multi-objective reward function, $R_{t}$, designed to balance fairness, throughput, and system stability at each time step $t$. The total reward is a summation of several components:
$R_{t}=R_{\text{fair}}+R_{\text{sum-rate}}+R_{\text{BER}}+R_{\text{power}}+R_{\text{stability}}$ (24)
Each component is defined as follows:
  • Fairness Reward ($R_{\text{fair}}$): To enforce equitable resource distribution, a significant reward is given for achieving near-perfect fairness, with a quadratic penalty for deviations. Jain’s fairness index, $J$, is calculated from the throughputs $T_{k}$ of $N$ users as defined in Eq. 25.
    $J=\frac{\left(\sum_{k=1}^{N}T_{k}\right)^{2}}{N\sum_{k=1}^{N}T_{k}^{2}}$ (25)
    The fairness reward is then calculated as:
    $R_{\text{fair}}=\begin{cases}+100,&\text{if }J\geq 0.99\\ -50\times(0.99-J)^{2},&\text{otherwise}\end{cases}$ (26)
  • Sum-Rate Bonus ($R_{\text{sum-rate}}$): To encourage overall performance, a bonus is given that is proportional to the throughput value in Mbps, which is obtained by dividing the total system throughput in bps (bits per second) by $10^{8}$.
    $R_{\text{sum-rate}}=w_{sr}\left(\frac{\sum_{k=1}^{N}T_{k}}{10^{8}}\right)$ (27)
    Here, $T_{k}$ represents the throughput of the $k$-th user in bps, and the weight coefficient $w_{sr}$ is set to 0.3.
  • BER Penalty ($R_{\text{BER}}$): A penalty is applied if the BER for any user exceeds a predefined threshold, ensuring communication reliability.
    $R_{\text{BER}}=\begin{cases}-20,&\text{if }\text{BER}_{k}>10^{-3}\text{ for any user }k\\ 0,&\text{otherwise}\end{cases}$ (28)
  • Power Efficiency Bonus ($R_{\text{power}}$): This component encourages the agent to select power allocation coefficients ($\rho$) that are not at the extremes, promoting a balanced allocation.
    $R_{\text{power}}=10\times(1-|\rho^{(1)}-0.25|)$ (29)
  • Stability Bonus ($R_{\text{stability}}$): A small reward is given for operating within a stable power allocation range, penalizing selections at the absolute boundaries.
    $R_{\text{stability}}=\begin{cases}+5,&\text{if }0.01\leq\rho^{(1)}\leq 0.49\\ -10,&\text{otherwise}\end{cases}$ (30)
An episode terminates successfully if the fairness target is met, yielding an additional bonus of +50. Conversely, failure conditions result in a penalty of -100.
@@
The SAC agent is trained for 1000 episodes by randomly generating user channel ratios and channels. Training is completed when the average reward over the last 100 episodes is 163. The agent is then tested with 1000 randomly generated pairs of $(h^{(1)},h^{(2)})$ where the value of $r$ is varied in increments of 0.001 within the range of 0 to 1.
@@
The changes in sum-rate, throughput, fairness, and BER values are depicted in Fig. 2(a), Fig. 2(b), Fig. 2(c) and Fig. 2(d) respectively, as the channel ratio $r$ varies. These values were obtained by applying proposed power allocation techniques to users in randomly generated $(h^{(1)},h^{(2)})$ channels. The performance curves for the SAC agent in the figures represent the average results obtained from 500 evaluation runs with different random seeds, confirming the stability and reliability of the learned policy. The figures also include the performance results of TDMA-based multiple access, GRPA, and NGDPA. Upon examination, it becomes clear that the trained agent demonstrates significantly superior performance in sum-rate, throughput, and fairness compared to other methods.
Algorithm 1 SAC Agent Training for NOMA-CSK Power Allocation
1:Input: State space $\mathcal{S}$, Action space $\mathcal{A}$; parameters $\gamma,\alpha,\tau,\eta$; buffer size $B$; batch size $M$.
2:Initialize: Actor network $\pi_{\phi}$, two critic networks $Q_{\theta_{1}},Q_{\theta_{2}}$, and target networks $Q^{\prime}_{\theta^{\prime}_{1}},Q^{\prime}_{\theta^{\prime}_{2}}$.
3:Initialize: Replay buffer $D$ with capacity $B$.
4:for each episode do
5:  Generate random channel gains $h^{(1)},h^{(2)}$ and get initial state $s$. $\triangleright$ State $s$ from Eq. 20
6:  for each step of the episode do
7:   Select action $a_{t}\sim\pi_{\phi}(a|s_{t})$. $\triangleright$ Action $a_{t}=\rho^{(1)}$ from Eq. 21
8:   Execute action $a_{t}$, calculate throughputs and BER.
9:   Observe reward $r_{t}$ and next state $s_{t+1}$. $\triangleright$ Reward $r_{t}$ from Eq. 24
10:   Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $D$.
11:   Sample a random mini-batch of $M$ transitions $(s_{i},a_{i},r_{i},s^{\prime}_{i})$ from $D$.
12:   $\triangleright$ Compute the target for the Q-functions
13:   With next actions $a^{\prime}_{i}\sim\pi_{\phi}(a^{\prime}|s^{\prime}_{i})$, the target is:
14:   $y_{i}=r_{i}+\gamma\left(\min_{j=1,2}Q^{\prime}_{\theta^{\prime}_{j}}(s^{\prime}_{i},a^{\prime}_{i})-\alpha\log\pi_{\phi}(a^{\prime}_{i}|s^{\prime}_{i})\right)$.
15:   $\triangleright$ Update the critic networks
16:   Update $\theta_{j}$ by minimizing the loss: $\mathcal{L}(\theta_{j})=\frac{1}{M}\sum_{i}(Q_{\theta_{j}}(s_{i},a_{i})-y_{i})^{2}$ for $j=1,2$.
17:   $\triangleright$ Update the actor network
18:   With new actions $\tilde{a}_{i}\sim\pi_{\phi}(a|s_{i})$, update $\phi$ by maximizing the objective:
19:   $J(\phi)=\frac{1}{M}\sum_{i}\left(\min_{j=1,2}Q_{\theta_{j}}(s_{i},\tilde{a}_{i})-\alpha\log\pi_{\phi}(\tilde{a}_{i}|s_{i})\right)$.
20:   $\triangleright$ Update the target networks
21:   Soft update target critic weights: $\theta^{\prime}_{j}\leftarrow\tau\theta_{j}+(1-\tau)\theta^{\prime}_{j}$ for $j=1,2$.
22:  end for
23:end for
@@
Figure 2: Performance comparison of the proposed SAC agent against GRPA, NGDPA, and TDMA in the trained environment.
For a 10 dB SNR, a BER of around $10^{-2}$ was achieved for the SAC agent, while the BER graphs obtained for GRPA and NGDPA are shown in Fig. 2(d). While not error-free at this low SNR, the agent’s power allocation is significantly more effective at minimizing errors compared to GRPA and NGDPA. The agent’s strategy ensures that the interference levels are managed well enough for the SIC process at the receiver to be highly successful relative to the alternatives, leading to superior communication reliability under the tested conditions. In contrast, the higher BER of GRPA and NGDPA indicates that their static allocation rules can lead to residual interference levels that are too high for reliable decoding.
As shown in Fig. 2(c), our SAC-based agent maintains a perfect fairness index across the entire range of channel gain ratios. In stark contrast, traditional schemes such as NGDPA suffer significant fairness degradation when users’ channel conditions become disparate, and GRPA performance deteriorates as the channel gains become more similar, while NGDPA’s fairness improves but at the cost of total throughput. These findings underscore the primary advantage of our adaptive SAC-based approach in ensuring equitable service across diverse channel environments. This consistent fairness is a direct outcome of the fairness centric sophisticated reward function design, which implements a quadratic penalty for fairness deviations and heavily rewards achieving the target fairness threshold. The agent effectively learns to balance maximizing the overall throughput, shown in Fig. 2(b), with the critical requirement of not starving any single user, which is a common pitfall for purely throughput-driven algorithms. It achieves the fairness level of an orthogonal method like TDMA while providing the superior spectral efficiency of NOMA.
To provide a concrete numerical illustration of these differences, Table 4 compares the performance of the SAC agent, GRPA, and NGDPA for a specific scenario where the channel gain ratio $r=0.5$. For this example, we selected a mid-range channel gain for the stronger user, $h^{(1)}=3.132\times 10^{-4}$, which results in $h^{(2)}=1.566\times 10^{-4}$. The table clearly shows the resulting power allocations, user throughputs, and BER performance, highlighting the practical implications of each strategy.
Table 4: Performance comparison for a specific scenario ($r=0.5$, $h^{(1)}=3.132\times 10^{-4}$)
Method
Power Alloc.
($\rho^{(1)}$)
Tput $U_{1}$
Tput $U_{2}$
BER $U_{1}$ BER $U_{2}$
SAC Agent 0.0333 102.03 102.21 1.52e-02 9.16e-03
GRPA 0.2000 100.36 82.93 3.14e-02 2.44e-02
NGDPA 0.3333 93.19 52.96 1.01e-01 9.23e-02
4.1 Computational Complexity Analysis
A crucial aspect for the practical deployment of any resource allocation scheme is its computational complexity. While the proposed SAC-based approach involves a training phase, this is a one-time, offline cost. The relevant metric for real-time performance is the inference complexity, which is the time taken to make a power allocation decision during operation.
Table 5 provides a qualitative comparison of the computational complexity of the SAC agent during inference against the conventional GRPA and NGDPA methods.
Table 5: Computational complexity comparison of power allocation schemes
Method Operation Type Real-time Complexity
GRPA/NGDPA
Direct calculation from
channel gain values
Low ($O(1)$)
SAC Agent
Forward pass through a
pre-trained neural network
Low, depends on network
size but constant per decision
As shown in the table, GRPA and NGDPA involve simple analytical calculations with constant low complexity. The SAC agent’s decision-making relies on a forward pass through its actor network. Although this is computationally more intensive than a simple formula, modern hardware can perform this operation in sub-millisecond timescales for the network sizes typically used in such applications. Therefore, the marginal increase in computational overhead is a justifiable trade-off for the significant improvements in system performance, fairness, and robustness demonstrated in our simulations. Essentially, we trade a negligible, constant-time inference cost for a dynamic, intelligent allocation policy that adapts to any channel condition, a capability that fixed-rule methods inherently lack.
4.2 Illumination Analysis
In order to assess the illumination performance of both the proposed and reference systems, we model the LED’s output PSD in two scenarios: without communication (baseline PSD) and under CSK modulation (altered PSD). The modulation process perturbs the average spectral distribution of the emitted light, thereby impacting the illuminance. The power contributions of the transmitted symbols for users $U_{1}$ and $U_{2}$ in color channel $i$ are given by Eq. 31. Using these contributions, the modified PSD, denoted $PSD_{a}(\lambda)$, is computed via Eq. 32, and this altered spectrum is subsequently employed to evaluate color metrics CRI, CCT, and luminous flux.
$TSP_{i}=\mathbb{E}\left(\left[S_{i}^{(1)}{\rho^{(1)}}+S_{i}^{(2)}{(1-\rho^{(1)})}\right]^{2}\right)$ (31)
$PSD_{a}(\lambda)=\sum_{i=1}^{3}PSD_{i}(\lambda)TSP_{i}$ (32)
In this case, after obtaining $PSD_{a}$, CRI, CCT and luminous flux parameters for the cases with and without communication are calculated using the open source program [32] and given in Table 6. The NOMA method did not change the CRI and CCT as expected. Only a decrease in luminous flux was observed due to modulation in communication. The stability of the CRI and CCT is a critical result for practical applications, as it ensures that the primary function of lighting is not compromised by the secondary data transmission task. The reduction in luminous flux is an anticipated and fundamental trade-off in VLC, as a portion of the optical power is utilized to carry data rather than solely for illumination.
Table 6: Illuminations parameters comparison table
CRI CCT Luminous flux (lm)
W/O Comm. 29.87 10105 5713
W Comm. CSK Only 29.84 10105 1714
W Comm. NOMA CSK 29.84 10105 1714
4.3 SNR-BER Analysis
To further analyze the performance under different conditions, we evaluated the BER as a function of SNR for three distinct channel ratios ($h^{(2)}/h^{(1)}\in\{0.1,0.5,0.9\}$). For this example, we selected a mid-range channel gain for the stronger user, $h^{(1)}=3.132\times 10^{-4}$ as before. The results are presented in Fig. 3. Across all evaluated scenarios, the proposed SAC agent consistently outperforms both GRPA and NGDPA, achieving lower BER for any given SNR. Notably, under high channel gain similarity (e.g., $r=0.9$), where user conditions are nearly uniform, GRPA exhibits the poorest performance, with the BER plateauing around $10^{-1}$ at 30 dB. Conversely, at lower channel gain ratios, NGDPA demonstrates the worst BER, rendering it impractical for reliable communication. This comprehensive analysis underscores the robustness of the SAC agent’s learned policy, which maintains superior performance across diverse channel conditions and a broad range of SNR levels.
(a) SNR vs. BER at $h_{2}/h_{1}=0.10$
(b) SNR vs. BER at $h_{2}/h_{1}=0.50$
(c) SNR vs. BER at $h_{2}/h_{1}=0.90$
Figure 3: BER performance comparison versus SNR for different channel ratios.